OpenStudy (anonymous):

solve the system of equation; (1/x)+(3/y)=(3/4) (3/x)-(2/y)=(5/12) Hint: m=(1/x) n=(1/y)

myininaya (myininaya):

m+3n=3/4 ....equation 1 3m-2n=5/12 .... equation 2 solve for m and n first then we will solve for x and y first step multiply equation 1 by (-3) then add this to equation 2

myininaya (myininaya):

let me know what you get after that step

OpenStudy (anonymous):

Thanks. I will.

OpenStudy (anonymous):

Wouldn't it be like by 4 to get rid of the denominator (3/4) ? Jusst a question.

myininaya (myininaya):

-3 (m +3n=3/4) -3m+3m-9n-2n=-9/4+5/12 -11n=-27/12+5/12 -11n=-22/12 -11n=-11/6 n=-11/6 (-1/11) n=1/6

OpenStudy (anonymous):

Antarctica comprises about 10.5% of the total land area on Earth. Antarctica has a surface area of about 5,400,000 square miles. What is the total land area on Earth?